OpenStudy (anonymous):

Which is 10^–4 written in standard form?

OpenStudy (anonymous):

10^(-4) = 1/10^4 = ???

OpenStudy (imstuck):

10^-4 means that the number is really small. Like 10^6, for example, would be a 1 with 6 zeros after it. This type of exponent makes the number very small. 1/10^4. The negative exponent requires you to move the constant to the denominator with a 1 in the numerator. 10^4 is a 1 with 4 zeros after it, so it is 1/10,000. One-ten thousandth.